OpenStudy (leozap1):

URGENT!!!! (FAN & MEDAL) PLEASE HELP. Which of the following equations would graph a line parallel to 3x-y=7

OpenStudy (leozap1):

y = 7x - 3 y = -3x + 7 y = -7 + 3 y = 3x

OpenStudy (nikato):

So 3x-y=7 Can you change it to slope intercept form? Hint: it'll look like y=mx+b where m=slope

OpenStudy (leozap1):

I think it might be y = 7x - 3

OpenStudy (leozap1):

I know it can't be y = -3x + 7, because it intersects.

OpenStudy (leozap1):

No its not y = 7x - 3

OpenStudy (leozap1):

The answer is y = 3x I had to work them ALL out. :(

OpenStudy (nikato):

Yes. D is correct but you didn't have to work through all that. So you have 3x-y=7 Add y to both sides. 3x=7+y Subtract 7 on both sides. 3x-7=y So the slope is 3. And you know parallel lines have the same slope So find the equation also with a slope of 3 which is D.y=3x